OpenStudy (anonymous):

the cost of a 25-cent candy bar is raised 20 percent. What is the new cost of the candy bar?

OpenStudy (zarkon):

25*1.20=30 cents

OpenStudy (anonymous):

.25+(.25*0.2)=.25+0.05 = 0.30 CENTS

OpenStudy (anonymous):

30 cents